OpenStudy (anonymous):

If 8.0 x 10-18 J of work is done to increase the electric potential difference of a test charge by 50.0 V, what is the magnitude of the charge? What formula do I use to prove that the answer is 1.6 x 10-19 C

OpenStudy (matt101):

Work done in an electric field is given by W=qΔV, where W is the work, q is the charge, and V is the voltage. Sub in the values given in the question and solve for q to get your answer. This equation is what it is because it's actually the same as the equation for work you're used to: W=Fd. In this case, we're dealing with an electrical force, which is given by F=kQq/d^2. Sub this into the work equation to find that W=kQq/d. Electric potential due to a point charge is given by V=kq/d, meaning you can rewrite the previous equation as W=qV to make it simpler.

OpenStudy (anonymous):

Thanks so much!